Upon the heating of black body the maximum of its spectrum shiftedfrom 690 nm to 500 nm. By what factor did the radiant emittance increase?

1
Expert's answer
2022-06-06T08:53:59-0400
"\lambda_{\max}=\frac{b}{T},\quad u=\sigma T^4"

"\frac{u_2}{u_1}=\frac{T_2^4}{T_1^4}=\frac{\lambda_{\max 1}^4}{\lambda_{\max 2}^4}=\frac{690^4}{500^4}=3.6"